Find the conversion factors of:
1. nm/min to μm/h
2. g/cn³ to kg/m³
1
Expert's answer
2020-05-11T20:04:21-0400

1) We know that "1\,\mu m = 10^{-6}\, m, \; \; 1\,nm = 10^{-9}\,m," so "1\,nm = \dfrac{10^{-9}}{10^{-6}}\,\mu m = 10^{-3}\, \mu m."

Next, we know that 1 hour is equal to 60 min. Therefore,

"1\dfrac{nm}{min} = 1 \dfrac{nm}{\dfrac{1}{60}h} = 60 \dfrac{nm}{h}." Then, "60 \dfrac{nm}{h} = 60 \dfrac{10^{-3}\,\mu m}{h} = 6\cdot10^{-2}\,\dfrac{\mu m}{h}."

If we have, for example, "v = 5 \, nm/min", it'll be equal to "5\cdot 6\cdot10^{-2}\,\mu m/h = 0.3\, \mu m/h."


2) We know that "1\,g = 10^{-3}\, kg" and "1\,cm^{3} = (10^{-2}\,m)^3 = 10^{-6}\,m^3." Therefore,

"1\dfrac{g}{cm^3} = 1\dfrac{10^{-3}\,kg}{10^{-6}\,m^3} = 10^3\dfrac{kg}{m^3}."

If we have, for example, "\rho = 4\,g/cm^3," then "\rho = 4\cdot10^3\,kg/m^3 = 4000\,kg/m^3."